40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 146 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  9601   Wed Feb 5 15:26:57 2014 manasaUpdateGeneralReady for pump down??

Quote:

 

 This sounds great! The only suggestion that I have is for checking POP. If you have the beam on the camera, you can hold a card in front of each mirror, and find out where the edge of the beam is. Introduce the card from the side, and watch for the point where you just start to see the beam on the camera be obstructed. Repeat for the other side, and you have an idea of the centering of the beam. 

I think this is most important for the in-vac mirrors, since the beam is large-ish, and we have to hit both steering mirrors at ~45 degrees.

 [EricQ, Manasa]

This check was done and we had to move one of the steering mirrors in pitch. Else, everything was just fine.

In-vacuum pictures of PR2 and PR3 new positions were taken. MC spot positions measured to be < 1mm and oplevs were centered.

  9600   Wed Feb 5 09:28:32 2014 SteveUpdateSUSETMY damping restored

ETMY damping restored.

  9599   Wed Feb 5 08:14:59 2014 SteveUpdateVAC vacuum computers are back without safety

Quote:

 

 1, Reset works

 2, Default values are lost. They actually reversed into open and turn on everything when power recycle c1vac1, c1vac2 or 24V dc power supply. This can vent the IFO in an event of power failure!

      It's may be the time to go back to an isolated, vacuum controller only computer.

 

The present valve configuration at Atm1

a, all annuloses are vented

b, valve cables disconnected at: VC1 and 4 IP gate valves

c, the RGA is off and it is pumped by the Maglev through VM2

d, cryo pump is being roughed with TP3  It's outgassing rate was 25 mTorr / min

e, Roughing hose is disconnected

 Vacuum checks before pump down:

1, check intermittent contact issue in black relay box

2, reconnect cable to ion pumps and vent them to atm before pump down

3, repeat power shutdown-reboot at atmosphere

  9598   Tue Feb 4 23:01:24 2014 JenneUpdateGeneralReady for pump down??

 

 This sounds great! The only suggestion that I have is for checking POP. If you have the beam on the camera, you can hold a card in front of each mirror, and find out where the edge of the beam is. Introduce the card from the side, and watch for the point where you just start to see the beam on the camera be obstructed. Repeat for the other side, and you have an idea of the centering of the beam. 

I think this is most important for the in-vac mirrors, since the beam is large-ish, and we have to hit both steering mirrors at ~45 degrees.

  9597   Tue Feb 4 17:40:19 2014 manasaUpdateGeneralReady for pump down??

Quote:

* PRM suspension has not been behaving well. PRM is being kicked around every 5-10 seconds when the PRC is aligned (as seen on REFL camera). We are not sure where this is coming from. The first time we saw this happening was when we were trying to lock PRC at low power even before we took the heavy doors off. So we are pretty sure this is not caused by the foil cover on the OSEMs. We tried turning ON/OFF the oplev servo, turning ON/OFF the damping loops and also checked the connections in the feedthrough and satellite box for the PRM. The OSEM sensor values for the suspension also seem to match the ones on the wiki.

 

This is solved.

The ASC for PRC for left turned ON. Turning it OFF solved the problem.

If there is no feedback regarding the POP alignment or anything to check with modified PRC length, we will close tomorrow morning.

  9596   Tue Feb 4 16:43:50 2014 manasaUpdateGeneralIFO alignment update

[EricQ, Manasa]

We are close to the end of the vent except for a couple of issues.

* POP is not visible on the IR card. But we see POP flashes unclipped on the camera and also spikes in POP DC. So  we are assuming that the POP path hasn't gone far off. If anybody has suggestions for a better method to check this, we could give it a try.

* PRM suspension has not been behaving well. PRM is being kicked around every 5-10 seconds when the PRC is aligned (as seen on REFL camera). We are not sure where this is coming from. The first time we saw this happening was when we were trying to lock PRC at low power even before we took the heavy doors off. So we are pretty sure this is not caused by the foil cover on the OSEMs. We tried turning ON/OFF the oplev servo, turning ON/OFF the damping loops and also checked the connections in the feedthrough and satellite box for the PRM. The OSEM sensor values for the suspension also seem to match the ones on the wiki.

 

Closeup checklist

 

  • Center beam on all AS optics
  • GET CAMERA IMAGES OF EVERYTHING

    • We must get images right before closing, right after closing, etc.
  • Make sure REFL is clear
    • dither PRM, see motion on AP tables
  • Make sure AS is clear
    • dither BS/ITM, see motion on AP tables
  • Using IPANG/POS pick-off mirrors, center beams on:
    • IPPOS
    • IPANG (IPANG aligned to be low in pitch. The beam comes out of vacuum unclipped and hits the first steering mirror outside vacuum at its lower edge)
  • Check green alignment in the arms and make sure the transmitted green reaches the PSL table.
  • Check all OpLevs centered, in and out of vacuum

  • Close PSL shutter & green shutters at the ends

  • Check jam nuts
  9595   Tue Feb 4 01:02:03 2014 KojiUpdateGreen LockingETMX green power

Manasa, Steve: Please revisit the Xend oven temperature again.


I found that the X end SLOW control was left on for ~15days. The output of the filter had grown to ~2e7.

This yielded the laser temperature pulled with the maximum output of the DAC.

This was the cause of the power reduction of the X end SHG; phase matching condition was changes as the wavelength of the IR was changed.

Once the SLOW output was reset, the green REFL was reduced from 4000cnt to 1800cnt.

Attachment 1: Screenshot-Untitled_Window.png
Screenshot-Untitled_Window.png
  9594   Tue Feb 4 00:42:18 2014 KojiUpdateGeneralX arm aligned for IR/GR

The X arm was also aligned for the IR by hand and ASS. Also the X end green PZT was aligned to make the TEM00 mode reasonably locked.

What I did:

- Looked at the ITMXF camera. It seemed that the green beam was hitting the mirror.

- Went to the end. Looked at the X end green REFL. Tuned coarse alignment of the ETMX so that the beam was (retro-)reflected to the Faraday and the REFL PD.

- Looked at the ETMX face from the view port. Tried to locate the spot from the ITMX by shaking the ITMX alignment with 0.1 and then 0.01 increments.

- After some struggle with the ETMX and ITMX alignment, resonant fringes were found on the ETMY face while I still looked at the ETMX.

- Once the ITMX/ETMX were aligned, the BS needed to be aligned. But of course there was no IR fringe.

- Returned to the original alignment of the ITMX to find the ITMX spot on the AS camera.
Then gradually moved the ITMX to the aligned value for the green while tracking the michelson alignment with the BS.
This made the AS spots at the upper left edge of the AS video image.

- This was enough to find the IR spikes at TRX. Then the ETMX was touched to maximize the transmission.

- Lock the cavity. Use the ASS to optimize the alignement.

- Once the arm mirrors were aligned, the Xend PZT was also adjusted to have TEM00 for the green beam.


Now I leave the IFO with ITMX/Y, ETMX/Y and BS aligned. As I wrote above, the AS spot is very high at the AS camera.
We need to revisit the AS steering (SR TTs?) to ensure the AS beam unclipped.

  9593   Mon Feb 3 23:31:33 2014 ManasaUpdateGeneralAlignment update / Y arm locked

[EricQ, Manasa, Koji]

We measured the spot positions on the MC mirrors and redid the MC alignment by only touching the MC mirror sliders. Now all the MC spots are <1mm away from the center.

We opened the ITMY and ETMY chambers to align the green to the arm. The green was already centered on the ITMY. We went back and forth to recenter the green on the ETMY and ITMY (This was done by moving the test masses in pitch and yaw only without touching the green pointing) until we saw green flashes in higher order modes. At this point we found the IR was also centered on the ETMY and a little low in pitch on ITMY. But we could see IR flashes on the ITMYF camera. We put back the light doors and did the rest of the alignment using the pitch and yaw sliders.

When the flashes were as high as 0.05, we started seeing small lock stretches. Playing around with the gain and tweaking the alignment, we could lock the Y arm in TEM00 for IR and also run the ASS. The green also locked to the arm in 00 mode at this point. We aligned the BS to get a good AS view on the camera. ITMX was tweaked to get good michelson.

  9592   Mon Feb 3 15:57:52 2014 SteveUpdateGreen LockingETMX green power

There was 0.2 mW green at the X end.

The doubling oven temp was changed from 37.5 to 36 degrees C

Power at green shutter 3 mW  The alignment was not touched.

Attachment 1: ETMXgreen.jpg
ETMXgreen.jpg
Attachment 2: XgreenOven36C.jpg
XgreenOven36C.jpg
  9591   Mon Feb 3 10:17:14 2014 SteveUpdatePEMdusty surfaces

Please wet WIPE before opening chamber or optical table ! ! 

 with methanol soaked kimwipes.

The Met One particle counter is located on CES wall, just behind ITMX chamber. 

The numbers are not so bad, but have you ( ...a)  asked the IFO lately?

Attachment 1: AP-ISCT.jpg
AP-ISCT.jpg
Attachment 2: BSC-rim.jpg
BSC-rim.jpg
Attachment 3: ITMX-ISCT.jpg
ITMX-ISCT.jpg
Attachment 4: 12daysAtm.png
12daysAtm.png
  9590   Fri Jan 31 19:29:36 2014 GabrieleSummaryLSCPRC and SRC lengths

 Today we measured the missing distance to reconstruct SRC length.

I also changed the way the mirror positions are reconstructed. In total for PRC and SRC we took 13 measurements between different points. The script runs a global fit to these distances based on eight distances and four incidence angles on PR2, PR2, SR2 and SR3. The optimal values are those that minimize the maximum error of the 13 measurements with respect to the ones reconstructed on the base of the parameters. The new script is attached (sorry, the code is not the cleanest one I ever wrote...)

The reconstructed distances are:

Reconstructed lengths [mm]:
LX    = 6771
LY    = 6734
LPRC  = 6752
LX-LY = 37
LSX   = 5493
LSY   = 5456
LSRC  = 5474

The angles of incidence of the beam on the mirrors are very close to those coming from the CAD drawing (within 0.15 degrees):

Reconstructed angles [deg]:
aoi PR3 = 41.11 (CAD 41)
aoi PR2 = 1.48  (CAD 1.5)
aoi SR3 = 43.90 (CAD 44)
aoi SR2 = 5.64  (CAD 5.5)

The errors in the measured distances w.r.t. the reconstructed one are all smaller than 1.5 mm. This seems a good check of the global consistency of the measurement and of the reconstruction method.

NOTES: in the reconstruction, the BS is assumed to be exactly at 45 degrees; wedges are not considered.

 

 

Attachment 1: map_jan31st.pdf
map_jan31st.pdf
Attachment 2: survey_v3.zip
  9589   Fri Jan 31 18:41:25 2014 manasaUpdateGeneralIFO alignment update

[EricQ, Gabriele, Manasa]

We found we had lost the Y arm pointing from yesterday. We tried to recover the pointing for a couple of hours and finally decided to take the ETMY heavy door off.

The input beam was aligned to the Y arm. We also got AS and REFL out of vacuum and on the cameras.

We put back the light doors and tried to lock the arms, but did not succeed as yet.

Things to do:
1. Lock arms for IR
2. Realign POP path
3. Recenter all oplevs
4. Try to check the state of PRC after the length change
5. Take in-vacuum pictures

  9588   Thu Jan 30 19:00:25 2014 GabrieleSummaryLSCPRC length changed

 [Manasa, EricQ, Gabriele]

Today we changed the PRC length translating PR2 by 27 mm in the direction of the corner. After this movement we had to realign the PRC cavity to get the beam centered on PRM, PR2, PR3, BS (with apertures) and ITMY (with aperture). To realign we had to move a bit both PR2 and PR3. We could also see some flashes back from the ETMY . //Edit by Manasa : We could see the ETMY reflection close to the center of the ITMY but the arm is not aligned or flashing as yet//. 

After the realignment we measured again the PRC length with the same method of yesterday. We only had to change one of the length to measure, because it was no more accessible today. The new map is attached as well as the new script (the script contains also the SRC length estimation, with random numbers in it).

The new PRC length is 6753 mm, which is exactly our target!

 

The consistency checks are within 5 mm, which is not bad.

We also measured some distances to estimate the SRC length, but right now I'm a bit confused looking at the notes and it seems there is one missing distance (number 1 in the notes). We'll have to check it again tomorrow.

 

 

 

Attachment 1: map_jan30.pdf
map_jan30.pdf
Attachment 2: survey_v3_jan30.m
clear all


global sos_lx sos_ly sos_cx sos_cy tt_lx ...
       tt_ly tt_cx tt_cy sos_sx sos_sy sos_dy

%% Survey of the PRC+SRC lengths %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%% measured distances
d_MB2_MY  = 2114 + 27 + 9;
... 446 more lines ...
  9587   Thu Jan 30 11:59:03 2014 manasaUpdateCDSfb timing was off

Quote:

Since this morning, the fb's timing has been off.  Steve pointed it out to me earlier today, but I didn't have a chance to look at it until now. 

This was different from the more common problem of the mx stream needing to be restarted - that causes 3 red blocks per core, on all cores on a computer, but it doesn't have to be every computer.  This was only one red block per core in the CDS FE status screen, but it was on every core on every computer. 

The error message, when you click into the details of a single core, was 0x4000.  I elog searched for that, and found elog 6920, which says that this is a timing issue with the frame builder.  Since Jamie had already set things on nodus' config correctly, all I did was reconnect the fb to the ntp: 

fb$ sudo /etc/init.d/ntp-client restart

As in elog 6920, the daqd stopped, then restarted itself, and cleared the error message. It looks like everything is good again.

I suspect (without proof) that this may have to do with the campus network being down this morning, so the computers couldn't sync up with the outside world.

The above timing problem has been repeating (a couple of times this week so far). It does not seem to be related to the campus network.

The same solution was applied.

  9586   Wed Jan 29 21:01:03 2014 GabrieleSummaryLSCPRC length measurement analysis

 [Manasa, EricQ, Gabriele]

We managed to measure the PRC length using a procedure close to the one described in slog 9573.

We had to modify a bit the reference points, since some of them were not accessible. The distances between points into the BS chamber were measured using a ruler. The distances between points on different chambers were measured using the Leica measurement tool. In total we measured five distances, shown in green in the attached map.

We also measured three additional distances that we used to cross check the results. These are shown in the map in magenta.

The values of the optical lengths we measured are:

LX    = 6828.96 mm
LY    = 6791.74 mm
LPRC  = 6810.35 mm
LX-LY = 37.2196 mm  

The three reference distances are computed by the script and they match well the measured one, within half centimeter:

M32_MP1  = 117.929 mm   (measured = 119 mm)
MP2_MB3  = 242.221 mm   (measured = 249 mm)
M23_MX1p = 220.442 mm   (measured = 226 mm)

See the attached map to see what the names correspond to.

The nominal PRC length (the one that makes SB resonant without arms) can be computed from the IMC length and it is 6777 mm. So, the power recycling cavity is 33 mm too long w.r.t. the nominal length. This is in good agreement with the estimate we got with the SB splitting method (4cm).

According to the simulation in the wiki page the length we want to have the SB resonate when the arms are there is 6753 mm. So the cavity is 57 mm too long.

Attached the new version of the script used for the computation.

Attachment 1: map.pdf
map.pdf
Attachment 2: script.zip
  9585   Wed Jan 29 16:36:37 2014 KojiSummaryGeneralHigh power beam blasting of the aLIGO RFPD

[Rich, Jay, Koji]

We blasted the aLIGO RF PD with a 1W IR beam. We did not find any obvious damage.
Rich and Jay brought the PD back to Downs to find any deterioration of the performance with careful tests.

The power modulation setup is at the rejection side of the PBS in front of the laser source.
I checked the beams are nicely damped.
As they may come back here tomorrow, a power supply and a scope is still at the MC side of the PSL enclosure.

  9584   Tue Jan 28 23:32:12 2014 KojiUpdateGeneralX/Y arm locked with the IR beam

[Koji EricQ]

The both arms have been locked with IR and aligned by ASS.

The IFO was left with ITMX/Y, ETMX/Y, BS, and PRM aligned, and the PSL shutter closed.


YARM
SIGNAL PATH:
POY11I(+45dB)->YARM(G=+1.0)->ETMY
NORM: TRYx10
TRIG: TRY 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay

XARM
SIGNAL PATH:
POX11I(+45dB)->XARM(G=+4.0)->ETMX
NORM: TRXx10
TRIG: TRX 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay


For decent locks, it was necessary that the offset of the error signals are trimmed at the input filters
even after running LSCoffset.py script.

Once the cavities were aligned for the IR, we could see the green beams are also flashing.
The Y arm was actually locked with the green with a TEM00 mode

Attachment 1: good_alignment.png
good_alignment.png
  9583   Tue Jan 28 22:24:46 2014 ericq UpdateGeneralFurther Alignment

[Masasa, ericq]

Having no luck doing things remotely, we went into the ITMX chamber and roughly aligned the IR beam. Using the little sliding alignment target, we moved the BS to get the IR beam centered on ITMX, then moved ITMX to get good michelson fringes with ITMY. Using an IR card, found the retroflection and moved ETMX to make it overlap with the beam transmitted through the ITM. With the PRM flashing, X-arm cavity flashes could be seen. So, at that point, both the y-arm and x-arm were flashing low order modes. 

  9582   Tue Jan 28 16:26:40 2014 SteveUpdateVAC vacuum computers are back without safety

Quote:

[Steve Koji]

We pushed the reset button of c1vac1 and c1vac2 and the vacuum screen is back.

First, we pushed the reset button of c1vac1 and pushed the one on c1vac2.
This did not bring c1vac2 up. We pushed the reset of c1vac2 again and now everything of the vacuum screen is back.

 1, Reset works

 2, Default values are lost. They actually reversed into open and turn on everything when power recycle c1vac1, c1vac2 or 24V dc power supply. This can vent the IFO in an event of power failure!

      It's may be the time to go back to an isolated, vacuum controller only computer.

 

The present valve configuration at Atm1

a, all annuloses are vented

b, valve cables disconnected at: VC1 and 4 IP gate valves

c, the RGA is off and it is pumped by the Maglev through VM2

d, cryo pump is being roughed with TP3  It's outgassing rate was 25 mTorr / min

e, Roughing hose is disconnected

Attachment 1: c1vac1_vac2areback.png
c1vac1_vac2areback.png
  9581   Tue Jan 28 11:13:50 2014 KojiUpdateVAC vacuum monitor is still blank

[Steve Koji]

We pushed the reset button of c1vac1 and c1vac2 and the vacuum screen is back.

First, we pushed the reset button of c1vac1 and pushed the one on c1vac2.
This did not bring c1vac2 up. We pushed the reset of c1vac2 again and now everything of the vacuum screen is back.

  9580   Tue Jan 28 09:51:47 2014 SteveUpdateIOOlow power pointing

 

PSL output is stable.

Attachment 1: IOO4dlowpr.png
IOO4dlowpr.png
  9579   Mon Jan 27 21:36:35 2014 ericq UpdateGeneralVent so far

After turning the slow FSS threshold down, the mode cleaner stays locked enough to do other things. We were able to align the tip tilts to the y-arm such that we were able to get some flashes in what looks like a TM00-ish mode. (It was necessary to align the PRM such that there was some extra power circulating in the PRC to be able to see the IR flashes on the ITMY face camera) This is enough to convince us that we are at least near a reasonable alignment, even though we couldn't lock to the mode. 

The x-arm was in a hairier situation; since the green beam wouldn't flash into any modes, we don't even know that a good cavity axis exists. So, I used the green input PZTs to shine the green beam directly on the earthquake stops on the ITMX cage, and then inferred the PZT coordinates that would place the green beam roughly on the center of ITMX. I moved the ETMX face camera such that it points at the ETMX baffle. I tried looking for the retroreflected green spot to no avail. Hopefully tomorrow, we can get ourselves to a reasonably aligned state, so we can begin measuring the macroscopic PRC length. 

  9578   Mon Jan 27 17:49:46 2014 ranaHowToComputer Scripts / Programssendmail started on nodus: fixing SwiftMail on Dokuwiki

Since the recent filesystem fracas, the new accounts could not be created on nodus / dokuwiki (for the controls workshop, for example).

I started sendmail on nodus using the command:   sudo /etc/init.d/sendmail start

and the SwiftMail plugin on there is now sending out the confirmation emails again. This will happen each time we reboot nodus, so let's replace it.

  9577   Mon Jan 27 12:26:00 2014 KojiUpdateIOOIOO Slow Actuator Servo threshold changed

In order to activate the slow actuator servo for the MC locking,
the threshold level for this servo (C1:PSL-FSS_LOCKEDLEVEL) was changed from 10000 to 700.

Now the servo started to move the PZT fast out to be controlled to 5V.

  9576   Mon Jan 27 09:08:00 2014 SteveUpdateVAC4 days at atmosphere

 

 

Attachment 1: day4atm.png
day4atm.png
Attachment 2: vent76bgTp3d4.png
vent76bgTp3d4.png
  9575   Sat Jan 25 21:09:16 2014 gabrieleHowToLSCProcedure to measure PRC length

I know the drawing is wrong. I put random distances, not realistic ones, and I did not try to get something close to reality. Once we put the measured distances, the drawing should (hopefully) be correct.

  9574   Fri Jan 24 13:10:12 2014 JamieHowToLSCProcedure to measure PRC length

Quote:

I wrote a MATLAB script that takes as input the measured distances and produce the optical path lengths. The script also produce a drawing of the setup as reconstructed, showing the measurement points, the mirrors, the reference base plates,  and the beam path. Here is an example output, that can be used to understand which are the five distances to be measured. I used dummy measured distances to produce it.

map.pdf

This path does not look correct to me.  Maybe it's because this is supposed to represent "optical path lengths" as opposed to actual physical location of optics, but I think locations should be checked.  For instance, PRM looks like it's floating in mid-air between the BS and ITMX chambers, and PR2 is not located behind ITMX.  Actually, come to think of it, it might just be that ITMX (or the ITMs in general) is in the wrong place?

Here is a similar diagram I produced when building a Finesse model of the 40m, based on the CAD drawing that Manasa is maintaining:

path.pdf

  9573   Fri Jan 24 12:44:25 2014 GabrieleHowToLSCProcedure to measure PRC length

Here is how to measure the PRC length with a set of distance measurements in the optical setup. 

We need to take distance measurements between reference points on each mirror suspension. For the large ones (SOS) that are used for BS, PRM and ITMs, the reference points are the corners of the second rectangular base: not the one directly in contact with the optical bench (since the chamfers make difficult to define a clear corner), but the rectangular one just above it. For the small suspensions (TT) the points are directly the corners of the base plates.

From the mechanical drawings of the two kind of suspensions I got the distances between the mirror centers and the reference corners. The mirror is not centered in the base, so it is a good idea to cross check if the numbers are correct with some measurements on the dummy suspensions.

I assumed the dimensions of the mirrors, as well as the beam incidence angles are known and we don't need to measure them again. Small errors in the angles should have small impact on the results.

I wrote a MATLAB script that takes as input the measured distances and produce the optical path lengths. The script also produce a drawing of the setup as reconstructed, showing the measurement points, the mirrors, the reference base plates,  and the beam path. Here is an example output, that can be used to understand which are the five distances to be measured. I used dummy measured distances to produce it.

map.pdf

In red the beam path in vacuum and in magenta the beam path in the substrate. The mirrors are the blue rectangles inside the reference bases which are in black. The thick lines are the HR faces. The green points are the measurement points and the green lines the distances to be measured. The names on the measurement lines are those used in the MATLAB script. 

The MATLAB scripts are attached to this elog. The main file is survey_v2.m, which contains all the parameters and the measured values. Update it with the real numbers and run it to get the results, including the graphic output. The other files are auxiliary functions to create the graphics. I checked many times the code and the computations, but I can't be sure that there are no errors, since there's no way to check if the output is correct... The plot is produced in a way which is somehow independent from the computations, so if it makes sense this gives at least a self consistency test. 

Attachment 2: survey_v2.m
global sos_lx sos_ly sos_cx sos_cy tt_lx tt_ly tt_cx tt_cy

%% Survey of the PRC length

%% measured distances
d_MB2_MY  = 2000.0;
d_MB3_MX  = 2000.0;
d_MB1_M31 = 400.0;
d_M32_M21 = 3000.0;
d_M22_MP  = 2000.0;
... 210 more lines ...
Attachment 3: distance.m
function d = distance(c1, c2)
    d = sqrt(sum((c1-c2).^2));
end
Attachment 4: draw_beam.m
function draw_beam(c1, c2, color)
    plot( [c1(1), c2(1)], [c1(2), c2(2)], color, 'LineWidth', 2)
end
Attachment 5: draw_measurement.m
function draw_measurement(c1, c2, color, name)
    plot( [c1(1), c2(1)], [c1(2), c2(2)], color)
    text( (c1(1)+c2(1))/2, (c1(2)+c2(2))/2 + 20, name, ...
        'FontSize', 5, 'HorizontalAlignment', 'center', ...
         'VerticalAlignment', 'middle')
end
Attachment 6: draw_point.m
function draw_point(c)
    plot(c(1), c(2), 'go', 'LineWidth', 2, 'MarkerSize', 3);
end
Attachment 7: draw_sos.m
function draw_sos(C, angle)
    global sos_lx sos_ly sos_cx sos_cy tt_lx tt_ly tt_cx tt_cy

    c(:,1) = [-sos_lx/2, -sos_ly/2 + sos_cy-sos_ly/2]';
    c(:,2) = [-sos_lx/2, sos_ly/2 + sos_cy-sos_ly/2]';
    c(:,3) = [sos_lx/2, sos_ly/2 + sos_cy-sos_ly/2]';
    c(:,4) = [sos_lx/2, -sos_ly/2 + sos_cy-sos_ly/2]';
    c(:,5) = [-sos_lx/2, -sos_ly/2 + sos_cy-sos_ly/2]';
    
    m_lx = 25.4*2;
... 18 more lines ...
Attachment 8: draw_tt.m
function draw_tt(C, angle)
    global sos_lx sos_ly sos_cx sos_cy tt_lx tt_ly tt_cx tt_cy

    c(:,1) = [-tt_lx/2, -tt_ly/2 + tt_cy-tt_ly/2]';
    c(:,2) = [-tt_lx/2, tt_ly/2 + tt_cy-tt_ly/2]';
    c(:,3) = [tt_lx/2, tt_ly/2 + tt_cy-tt_ly/2]';
    c(:,4) = [tt_lx/2, -tt_ly/2 + tt_cy-tt_ly/2]';
    c(:,5) = [-tt_lx/2, -tt_ly/2 + tt_cy-tt_ly/2]';
    
    m_lx = 25.4;
... 18 more lines ...
  9572   Thu Jan 23 23:10:19 2014 ericq UpdateGeneralVent so far

[ericq, Manasa, Jenne]

Summary: We opened up the BS and both ITM chambers today, and put the light doors on. //Edit : Manasa  Post-vent the MC was very much misaligned in yaw. Both the ITMs moved in pitch as inferred from the oplev; but there is still light on the oplev PDs//. We toiled with the PMC and mode cleaner for a while to get reasonable transmission and stability (at least for a period of time). We then tried to lock IR to the y-arm, to no avail. 

Locking the PMC doesn't seem very robust with the low power level we have; adjusting the gain at all when it's locked throws it right out. The mode cleaner spot was visibly moving around on MC2 as well. We'll continue tomorrow. 

Details about alignment efforts: Manasa and I tried for a while to try and align the y-arm for IR. Straight out of venting the green TM00 would lock to the y-arm with about .45, as compared to .8 before venting, so it didn't seem to drift too far. The x-arm would even flash any modes, however. For a while, IR was no where to be seen after the mode cleaner. Eventually, we used the tip tilts to bring the AS beam onto the camera, which exhibited fringes, so we knew we were hitting the ITMs somewhere. We wandered around with the ETM to see if any retroflection was happening, and saw the IR beam scatter off of the earthquake stop. We moved it to the side to see it hitting the OSEM holder, and moved down to the bottom OSEM holder to get an idea of where to put pitch to get roughly the center of the ITM, then undid the yaw motion.

There, we would see very infrequent, weak flashes. We weren't able to distinguish the mode shape though; however, the flashes were coincident with where the green would lock to a very yaw-misaligned fishbone mode, to the lower right of the optic's center. We figured that if we gradually fixed the green alignment with the mode shapes we could see and actually lock on, we could use the tip tilts to adjust the IR pointing and keep it coincident and eventually resonate more. However, this didn't really work out. The flashes were very infrequent, and at this point the PMC/MC were getting very touchy, and would cease to stay locked for more than a minute or two. At this point, we stopped for the day. 

 

  9571   Thu Jan 23 13:23:10 2014 SteveUpdateVAC40m IFO is at atmosphere

Quote:

 

There is BLANK VacControl_BAK.adl screen only. 

I can move a valve by disconnecting it's  solenoid power if it's position is normally open.

I will close V1 and check computer cable connections and move on with manual - hand disconnect ea valve to be moved into the right position for vent. Valve positions will be confirmed by looking manual indicators on valves.

 The 40m vacuum envelope vent is completed with instrument grade air.

Valve configuration: chamber open, RGA is pumped through VM3 by TP3,

  9570   Thu Jan 23 08:04:04 2014 SteveUpdateVACvacuum control screen is blank

 

There is BLANK VacControl_BAK.adl screen only. 

I can move a valve by disconnecting it's  solenoid power if it's position is normally open.

I will close V1 and check computer cable connections and move on with manual - hand disconnect ea valve to be moved into the right position for vent. Valve positions will be confirmed by looking manual indicators on valves.

Attachment 1: pd76m170d2dRgaOn.png
pd76m170d2dRgaOn.png
  9568   Wed Jan 22 20:00:41 2014 JenneUpdateGeneralVENT GO!

Steve, please begin the vent!!

[EricQ, Jenne]

We have followed the pre-vent checklist, and done everything except check the jam nuts (which Steve can do in the morning).

We are ready to vent, so Steve, please begin bringing us up to atmosphere first thing in the morning.

Here is a copy of the list, from the wiki:

 

 

  • Center all oplevs/IPPOS/IPANG
  • Align the arm cavities for IR and align the green lasers to the arms. (Green powers were both ~0.8.  We only touched the Xend PZTs remotely, did not touch Yend).
  • Make a record of the MC pointing
  • Align the beam at the PSL angle and position QPDs (Did not need doing, left QPDs as-is so we keep our long-term trend.)
  • Reduce input power by touching wave plate on the PSL table BEFORE THE PMC.  (HWP was at 269degrees, now at 3 hundred something to make power just before PSL shutter 90mW)
  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  • Close shutter of PSL-IR and green shutters at the ends
  • Make sure the jam nuts are protecting bellows
  •  

     

    Attachment 1: IFOstatus_lowPower_preVent.png
    IFOstatus_lowPower_preVent.png
      9567   Wed Jan 22 18:17:46 2014 JenneUpdateCDSfb timing was off

    Since this morning, the fb's timing has been off.  Steve pointed it out to me earlier today, but I didn't have a chance to look at it until now. 

    This was different from the more common problem of the mx stream needing to be restarted - that causes 3 red blocks per core, on all cores on a computer, but it doesn't have to be every computer.  This was only one red block per core in the CDS FE status screen, but it was on every core on every computer. 

    The error message, when you click into the details of a single core, was 0x4000.  I elog searched for that, and found elog 6920, which says that this is a timing issue with the frame builder.  Since Jamie had already set things on nodus' config correctly, all I did was reconnect the fb to the ntp: 

    fb$ sudo /etc/init.d/ntp-client restart

    As in elog 6920, the daqd stopped, then restarted itself, and cleared the error message. It looks like everything is good again.

    I suspect (without proof) that this may have to do with the campus network being down this morning, so the computers couldn't sync up with the outside world.

      9566   Wed Jan 22 16:36:45 2014 ericqUpdateElectronicsRF distribution box power button fail

    Quote:

    Rana, Gabriele and I are trying to measure the FSR of the PRC (elog about that later), and we turned off the power to the RF generation box so that we could switch cables at the EOM combiner.  However, as in elog 9101, the power button won't latch when we try to turn the power back on.  All 3 of us tried, to no avail.  For our measurement, poor Gabriele is standing holding the button pushed in, so that we can have some RF sidebands. 

    Tomorrow, we'll have to pull the RF generation box, and put in a better switch.

    I replaced the stupid broken fancy button with a simple sturdy switch. I had to file out the hole in the chassis a bit, but the switch is pressed in tightly and securely. I put the box back in the rack, but the power cable was coming directly from the power supplies with no fuses. The box was drawing ~.9 and 1.5 Amps from two supplies, so I put 2A fuses on both. Plugged everything back in, and the mode cleaner locks, so it looks like all is well.

    RXA: When its so close, I prefer to size it up by 1 step. Please change to 5A fuses. Otherwise, we may blow them from power glitches.

    Q: 5A fuses have been swapped in

      9565   Wed Jan 22 15:24:11 2014 SteveSummaryVACRga scan after reboot

    Quote:

    [Jenne, Steve]

    Steve noticed that the RGA was not logging data and that not all the correct connection lights were on, and he wasn't able to run the "RGAset.py" script (in ...../scripts/RGA/) that sets up the proper parameters. 

    I looked, and the computer was not mounting the file system.  I did a remote shutdown, then Steve went in and pushed the power button to turn the machine back on.  After it booted up, it was able to talk to the file system, so I started ..../scripts/RGA/RGAset.py .  The first 2 times I ran the script, it reported errors, but the 3rd time, it reported no communication errors.  So, now that the computer can again talk to the file system, it should be able to run the cronjob, which is set to take data at 4am every day.  Steve will check in the morning to confirm that the data is there.  (The last data that's logged is 22Dec2013, 4am, which is right around when Koji reported and then fixed the file system).

     

     

     We are venting tomorrow. This give us an opportunity to fix sick vacuum controller computer. Jamie volunteered.

    Remember to rough down cryo and ion pump volumes. Their pressure can be at 1 Torr range after years of accumulated outgassing.  Without total valve controls it is dangerous to have these air pockets. Some of their  gate valves can be leaking and that would explane the slower pumpdown speed. 

     

    Attachment 1: Rgascan169dwarmingup.png
    Rgascan169dwarmingup.png
    Attachment 2: BlankMedmMustBeFixed.png
    BlankMedmMustBeFixed.png
      9564   Wed Jan 22 09:05:42 2014 GabrieleConfigurationLSCREFL11 back

     Yesterday night I plugged back the REFL11 RF cable into the corresponding demodulation board.

      9563   Tue Jan 21 19:41:59 2014 JenneUpdateElectronicsRF distribution box power button fail

    Rana, Gabriele and I are trying to measure the FSR of the PRC (elog about that later), and we turned off the power to the RF generation box so that we could switch cables at the EOM combiner.  However, as in elog 9101, the power button won't latch when we try to turn the power back on.  All 3 of us tried, to no avail.  For our measurement, poor Gabriele is standing holding the button pushed in, so that we can have some RF sidebands. 

    Tomorrow, we'll have to pull the RF generation box, and put in a better switch.

      9562   Tue Jan 21 17:26:59 2014 JenneSummaryVACRebooted RGA computer and reset RGA settings

    [Jenne, Steve]

    Steve noticed that the RGA was not logging data and that not all the correct connection lights were on, and he wasn't able to run the "RGAset.py" script (in ...../scripts/RGA/) that sets up the proper parameters. 

    I looked, and the computer was not mounting the file system.  I did a remote shutdown, then Steve went in and pushed the power button to turn the machine back on.  After it booted up, it was able to talk to the file system, so I started ..../scripts/RGA/RGAset.py .  The first 2 times I ran the script, it reported errors, but the 3rd time, it reported no communication errors.  So, now that the computer can again talk to the file system, it should be able to run the cronjob, which is set to take data at 4am every day.  Steve will check in the morning to confirm that the data is there.  (The last data that's logged is 22Dec2013, 4am, which is right around when Koji reported and then fixed the file system).

     

     

      9561   Fri Jan 17 11:44:25 2014 GabrieleSummaryLSCMore length measurements, more confusion

     I analyzed the data taken yesterday. 

    The AS11 data in PRMI configuration is very bad, while the AS55 seems good enough:

    results_as11.pdfresults_as55.pdf

    The phase differences are 

    AS11 = 21 +- 18 degrees (almost useless due to the large error)

    AS55 = 11.0 +- 0.4 degrees

    The AS55 phase difference is not the same measured in the last trial, but about half of it. The new length estimates are:

    AS11 = 3.2 +- 2.8 cm

    AS55 = 0.47 +- 0.01 cm

    We can probably forget about the AS11 measurement, but the AS55 result is different from the previous estimate... Maybe this is due to the fact that Eric adjusted the PRCL offset, but then we're going in the wrong direction....

      9560   Thu Jan 16 21:38:13 2014 ericqUpdateLSCRepeat of PRC length measurement

    [ericq,Jenne]

    Since we don't have agreement between the measurements we made the other day and the earlier estimations, I wanted to repeat the demodulation angle measurement. We had to do a few things to keep the PRMI locked, since in the last few days, it hasn't been stable enough.

    The mode cleaner had been very fussy lately; the WFS were pushing in a way that caused fast oscillations of the transmission and reflection powers. I turned off the servos, manually aligned the mode cleaner to transmission of about 15k and refl of about .4, centered the beams on the WFS QPDs, and turned the loops back on. Things were much stable after that. Also, Jenne noticed that the PMC loop had walked the laser PZT temperature to a bad place, and fixed it.

    After aligning the carrier locked PRMI, the last piece needed to get things stable enough for sideband locking was turning off the angular damping on the PRM suspension screen (this was turned back on when we were done). Waiting until evening noise levels probably helped too. We used a 1000 count MICH excitation in the PRMI case, and recorded data for about a minute in one degree steps around the demodulation phase that looked to put the excitation entirely within the Q of the PD. Also, we notched out the excitation frequency in the MICH servo bank for today's measurement; I think it's outside of the loop bandwidth anyways, but it's good to be sure. 

    Jenne and I pondered a bit whether changing the AS55 demodulation phase while it (AS55 Q) is being used as the MICH control signal introduces subtleties that we haven't anticipated, but couldn't come up with anything concrete. Changing the angle from the what maximizes the Q just looks like a slight change in MICH gain, and shouldn't affect the phase of the excitation signal on the PD...

    In any case, the data have been recorded, and the results will follow soon. 

      9559   Thu Jan 16 08:19:29 2014 SteveUpdatePEM4.4 and 3.8M local earth quakes


    It looks like there was a 4.4 magnitude earthquake near Fontana, CA around 1:30am today.  This tripped all of the suspension watchdogs, which Q has just now re-enabled. 

       Earth quake shake down yesterday Atm1

    Atm2, today's shake

    Attachment 1: local4.4eq.png
    local4.4eq.png
    Attachment 2: local3.8eq.png
    local3.8eq.png
      9558   Wed Jan 15 18:42:57 2014 JenneUpdateASCPOP ASC QPD offline for a few hours this afternoon

    I was in the lab, near the south end of the ITMX oplev table, looking for something, and I bumped the POP ASC QPD's power supply.  I thought that it was fine, but did not adequately check it.  When EricQ asked me just now about why the PRC is so wobbly today, I checked, and the power for the QPD wasn't properly connected (it's kind of a crappy connector, that if you nudge, contacts or loses contact).  Anyhow, I restored power to the QPD, and the PRC looks a little more stable now.  My fault for not checking more carefully, and my apologies to Q and Gabriele for their frustrations this afternoon.

      9557   Wed Jan 15 18:18:15 2014 GabrieleSummaryLSCPRC length measurement analysis

     I analyzed the data we took yesterday, both using AS11 and AS55. For each value of the phase I estimated the Q/P ratio using a demodulation code. Then I used a linear regression fit to estimate the zero crossing point.

    Here are the plots of the data points with the fits:

    experimental_data_as11.pdfexperimental_data_as55.pdf

    The measurements a re more noisy in the PRMI configuration, as expected since we had a lot of angular motion. Also, the AS11 data is more noisy. However, the estimated phase differences between PRMI and MICH configurations are:

    • AS11 = -10.9 +- 1.0 degrees
    • AS55 = -21.1 +- 0.4 degrees

    The simulation already described in slogs 9539 and 9541 provides the calibration in terms of PRC length. Here are the curves

    simulation_phase_difference.pdf

    The corresponding length errors are

    • AS11 = 1.44 +- 0.13 cm
    • AS55 = 0.59 +- 0.01 cm

    The two results are not consistent one with the other and they are both not consistent with the previous estimate of 4 cm based on the 55MHz sideband peak splitting.

    I don't know the reason for this incongruence. I checked the simulation, repeating it with Optickle and I got the same results. So I'm confident that the simulation is not completely wrong.

    I also tried to understand which parameters of the IFO can affect the result. The following ones have no impact

    • Beam matching
    • ITM curvatures
    • Schnupp asymmetry
    • Distance PR-BS
    • ITM and PRM misalignments

    The only parameters that could affect the curves are offsets in MICH and PRCL locking point. We should check if this is happening. A first quick look (with EricQ) seems to indicate that we indeed have an offset in PRCL. However, tonight the PRMI is not locking stably on the sidebands. 

    If possibile, we will repeat the measurement later on tonight, checking first the PRCL offset.

     

      9556   Wed Jan 15 13:05:30 2014 JenneUpdatePEM4.4 EQ in Fontana, CA

    It looks like there was a 4.4 magnitude earthquake near Fontana, CA around 1:30am today.  This tripped all of the suspension watchdogs, which Q has just now re-enabled. 

      9555   Tue Jan 14 19:10:51 2014 GabrieleSummaryLSCPRC length measurement

     [EricQ, Gabriele]

    We could carry out the measurement of PRC length. The AS110 photodiode was plugged into REFL11. So REFL11 is giving us the AS11 signal. Here is the procedure.

    1. Lock MICH.
    2. Add a line in MICH (amplitude 20000 counts)
    3. Tune AS11 demod phase to have the line in I.
    4. Change the demod phase by steps of 1 degree around the rough optimum, taking one minute of data at each point
    5. Lock PRMI on sidebands
    6. Add a line in MICH (amplitude 500 counts)
    7. Tune AS11 demod phase to have the line in I.
    8. Change the demod phase by steps of 1 degree around the rough optimum, taking one minute of data at each point

    We repeated the same measurement also using AS55, with the same procedure.

    Roughly, the phase difference for AS11 was 11 degrees and for AS55 it was 23 degrees. A more detailed analysis and a calibration in terms of PRC length will follow.

      9554   Tue Jan 14 19:05:48 2014 GabrieleSummaryLSCPRMI locked on sideband

    [Gabriele, EricQ] 

    Finally, we managed to lock PRMI on sidebands:

    • MICH locked on AS55_Q with gain -10. Demod phase of AS55=18
    • PRCL locked on REFL55_I with gain -0.04. Demod phase of REFL55=88
    • Triggering on POP110_I at 50/10
    • Filter "1:5" of MICH engaged, this improved a lot the stability

     

      9553   Tue Jan 14 10:34:57 2014 SteveUpdatePSLgreen transmission measurment

    GariLyn is using our green light on the west side of the PSL table. The green PDA36As were moved and the HEPA turned up to 60V

    Attachment 1: greenPickUp.jpg
    greenPickUp.jpg
      9552   Tue Jan 14 10:12:12 2014 SteveUpdatePSLlaser drift monitor set up idea

    Quote:

    this locationQuote:

    Quote:

    Quote:

    I wonder what's drifting between the laser and the PMC? And why is it getting worse lately?

     The PMC refl is bad in pitch today, and the transmission is only 0.76, rather than our usual 0.83ish.

    I did a quick, rough tweak-up of the alignment, and now we're at 0.825 in transmission.

     The PMC transmission continuously degrades. In order to see what is really drifting the laser output after PBS was sampled as shown.

     IOO pointing is drifting in pitch. I'd like to use a QPD instead of the paper target to see if the Innolite output is stable. The idea is to move temporarily IOO-QPD_POS to  this location

     I do like to move IOO-QPD_POS temporarily to see that the feedback has anything to do with with the pointing.

    Attachment 1: bad4thday.png
    bad4thday.png
      9551   Mon Jan 13 19:31:04 2014 GabrieleSummaryLSCPRMI locking

    [Gabriele, EricQ]

    We wanted to try the PRC length measurement,but we ended up spending all the afternoon to lock the PRMI on sidebands. Here are some results

    • Lock of PRM on carrier is easy with MICH on AS55_Q and PRCL on REFL55_I. The optimal gains to avoid correction gain peaking are MICH=-5 and PRCL=0.012. The alignment today was must more stable over time than on Friday
    • We wanted to move MICH on REFL55_Q. After a few trials we discovered that REFL55_Q is not seeing any MICH signal at all. This is quite strange and we don't understand this.
    • Locking PRMI on carrier using REFL165_I and Q was difficult. We thought that was due to a RF amplifier installed in December. We tried to remove it, but since this did not help, we put it back
    • We could lock PRMI on carrier and sidebands using REFL11 signals. The optimal demodulation phase is 155. Lock on carrier was achieved with gains MICH=5, PRCL=0.7. Lock on sidebands with MICH gain=-10 and PRCL gain=-0.3. The MICH correction is very high and exciting a tone at few hundreds Hz. Maybe a violin mode of the ITMs?
    • Finally we could tune the REFL165 phase to 118.5 and lock on carrier. The lock was not very stable. Gain are: lock on carrier MICH=-0.5, PRCL=0.05. The error signal has a very big offset: to get a dark fringe we had to add a MICH offset of 500 counts. We also had to engage the CLP400 filter to avoid saturating ITMs corrections with MICH.
    ELOG V3.1.3-